# SelaVPR++: Towards Seamless Adaptation of Foundation Models for Efficient Place Recognition

### v1

@@ IV-D Ablation Study
It is worth noting that the above fine-tuning methods are only applied to the backbone, while the aggregator (the GeM layer and the two linear layers before and after it) is always trainable. We use both DINOv2-base and DINOv2-large for experiments, outputting 2048-dim and 4096-dim global features for retrieval respectively. The GSV-Cities dataset is used for training. Since the Full-tuning will take up a huge amount of GPU memory, we adjust the batch size to 60 for all models based on DINOv2-large. Even so, the Full-tuning method still requires 6 NVIDIA 3090 GPUs. The results are shown in Table IV (DINOv2-base) and Table V (DINOv2-large). We summarize the two tables in the following points: (1) For the Freezing method, since our aggregator contains two trainable linear layers that can also adapt the features produced by backbone to the VPR task, it achieves decent results for partial datasets. However, for the Nordland dataset, which is mainly collected in natural scenes and across seasons, the model needs to be able to capture detailed information in the image that can distinguish places. Simply adjusting the feature map output by the frozen backbone cannot effectively bridge the gap between the tasks of model pre-training and VPR. Therefore, the performance of this method is much lower than other methods. (2) As for Full-tuning, when using the DINOv2-base backbone, the difference between it and other methods is not large on most datasets. However, it is obviously inferior to other methods on Tokyo247, which contains many nighttime images. It is because there is a generalization gap between Tokyo24/7 and the GSV-Cities training dataset (without night images), and full fine-tuning damages the excellent representation ability of the pre-trained foundation model (i.e., catastrophic forgetting). This phenomenon is more obvious for DINOv2-large that has a larger number of parameters. (3) The performance of the PEFT methods (i.e., SelaVPR-global, MemEfficient-vanilla, and MemEfficient-MCA) when using DINOv2-large is generally better than that of DINOv2-base. However, the R@1 results of the Full-tuning and Partial-tuning methods on Tokyo247 decrease when using DINOv2-large. Especially for Full-tuning, the R@1 on Pitts30k and MSLS-val also decreases. This shows that the PEFT method is more suitable for large models than direct (full or partial) fine-tuning. (4) Both the Partial-tuning and PEFT methods achieve good performance on all these datasets. However, compared with the previously proven Partial-tuning and SelaVPR, the memory-efficient adaptation method used in this paper can achieve better results. Furthermore, our MemEfficent-MCA upgrades vanilla adapters to MultiConv adapters, which can facilitate feature interactions along the spatial axes and introduce the multi-scale local priors. Thus, it achieves the best performance among these methods.
@@
TABLE IV: The Results of Different Fine-Tuning Method with the DINOv2-Base Backbone. Models Are Fine-Tuned on GSV-Cities and Batch Size Is Set to 120.
@@
TABLE V: The Results of Different Fine-Tuning Method with the DINOv2-Large Backbone. Models Are Fine-Tuned on GSV-Cities and Batch Size Is Set to 60.
@@
TABLE VI: The Comparison of Different Fine-Tuning Methods in Training Time for a Single Epoch on GSV-Cities. MemEfficient-MCA(4) Indicates Only Applying 4 Adapters for the Last 4 Blocks. We Consistently Train the Models on an NVIDIA A800 GPU with Batch Size Set to 40.
@@
Meanwhile, since the gradient backpropagation in our method no longer passes through the backbone during training, it saves a lot of training time. Table VI shows the training time of different fine-tuning methods for training an epoch on GSV-Cities (using an NVIDIA A800). Compared with Full-tuning, SelaVPR-global does not save significant time, that is, it is also not efficient in training time. For DINOv2-base, our MemEfficient-MCA consumes only 6.85 min/epoch, which is less than Partial-tuning (both are only slightly more than Freezing). For DINOv2-large, our method applies adapters to the last 16 blocks (DINOv2-base has only 12 blocks), while Partial-tuning still only fine-tunes the last 4 blocks. Therefore, the latter achieves a slightly shorter training time. However, our approach can further improve efficiency by reducing the number of adapters. When we use only 4 adapters, i.e., MemEfficient-MCA(4), the training time is 16.3% less than Partial-tuning. These results show that our method is efficient in training time in addition to parameter and memory. In the next ablation experiment, we will demonstrate that our method can still achieve excellent recognition performance when using only a small number of adapters.
TABLE VII: The Results of Using Different Number of Adapters (with Two Ways). All Models Are Based on DINOv2-Large Backbone and Trained on Our Unified Dataset (Batch Size is 120).
@@
R@1 R@5 R@10 R@1 R@5 R@10 R@1 R@5 R@10 R@1 R@5 R@10 R@1 R@5 R@10
Freezing 92.7 96.8 97.7 92.0 96.8 97.6 95.6 98.4 99.0 69.7 82.8 87.3 87.5 93.7 95.4
All 94.3 97.4 98.0 94.7 97.7 98.1 97.5 98.7 99.4 96.2 98.8 99.4 95.7 98.2 98.7
Every 2 94.3 97.2 97.9 94.6 97.3 98.1 97.1 99.7 99.7 96.5 98.9 99.3 95.6 98.3 98.8
Every 3 94.4 97.5 98.1 94.2 97.3 97.8 97.5 98.7 98.7 96.4 98.9 99.5 95.6 98.1 98.5
Every 6 93.0 96.8 97.7 93.8 96.9 97.6 97.5 98.4 99.4 93.7 97.8 98.7 94.5 97.6 98.4
Last 16 94.4 97.5 98.1 94.3 98.0 98.2 98.1 99.0 99.7 97.0 99.1 99.5 96.0 98.4 98.9
Last 12 94.3 97.5 98.1 94.7 97.7 98.1 97.8 98.7 99.4 96.6 98.8 99.2 95.9 98.2 98.7
Last 8 94.0 96.8 97.7 94.9 97.6 98.0 97.1 99.7 100.0 95.5 98.6 99.1 95.4 98.2 98.7
Last 4 94.0 97.3 98.0 94.1 97.2 97.8 97.5 99.4 99.4 92.5 97.2 98.4 94.5 97.8 98.4
@@
2) Effect of the number of adapters. As described in Section III-B, our approach can further improve efficiency by reducing the number of adapters in two ways, namely, applying one adapter per $m$ blocks or applying adapters only to the last partial blocks. We equip DINOv2-large with different numbers of adapters and train them on our unified dataset, yielding the results shown in Table VII. All methods equipped with adapters achieve good results and are significantly better than the baseline (i.e., Freezing). We mainly observe the performance of each setting through the average of R@N, which shows that the more adapters used in both ways, the better the performance. An exception is that the performance of applying adapters to the last 16 blocks is better than that for all blocks, indicating that the former has basically reached saturation performance. Therefore, for DINOv2-large, our recommendation is to only apply adapters to the last 16 blocks. In addition, the number of adapters used when applying an adapter per 2, 3, and 6 blocks just corresponds to applying adapters to the last 12, 8, and 4 blocks, and the results of the two are very close. So we believe both ways are good choices. Finally, reducing the number of adapters for the sake of efficiency is also advisable, because our method with only 4 adapters also achieves promising performance.
3) Effect of our novel two-stage paradigm. Our novel two-stage VPR paradigm discards the local features used in SelaVPR, and instead only uses global features for both initial retrieval and re-ranking. To demonstrate our re-ranking paradigm, we compare the performance of direct retrieval using binary features and floating-point features, as well as our re-ranking method (binary features for initial retrieval and floating-point features for re-ranking). We conduct experiments using both the DINOv2-base and DINOv2-large backbones with our unified dataset for training. The results are shown in Table VIII. Since the binary features we use are also low-dimensional, their performance (R@1/R@5/R@10) is significantly lower than that of high-dimensional floating-point features. For the model using DINOv2-base, the R@1 margins on Pitts30k, MSLS-val, Tokyo247, and Nordland are 4.3%, 4.2%, 6.3%, and 16.3%, respectively. However, in our re-ranking pipeline, the initial search (using binary features) is only expected to return the top-100 candidate images with correct results. The difference between the R@100 results obtained by binary and floating-point features is very slight. After re-ranking the top-100 candidates using high-dimensional floating-point features, the R@1/R@5/R@10 results of our method are not significantly different from those yielded by directly using high-dimensional floating-point features (some are slightly higher, and some are slightly lower).
Since our method uses compact binary features to provide candidate images by calculating Hamming distances, this can further greatly improve retrieval efficiency. Table IX shows the retrieval latency (including initial retrieval latency and re-ranking latency) of a single query on Pitts30k for three two-stage methods, i.e., TransVPR [12], SelaVPR [13], and SelaVPR++. We also provide the latency of direct retrieval with the L2 distance for 4096-dim, 2048-dim, and 512-dim floating-point features, as well as with the Hamming distance for 512-dim binary features. Although our previous work SelaVPR matches local features for re-ranking, it eliminates the time-consuming geometric verification process in the common two-stage method (e.g., TransVPR). Its re-ranking latency is thus more than one order of magnitude lower than that of TransVPR. Our SelaVPR++ directly uses global features for re-ranking, which is further more than two orders of magnitude faster than SelaVPR. As for the initial retrieval latency, once the measurement is determined (e.g., L2 distance), the latency is roughly proportional to the feature dimension. When the feature dimension is 4096, the retrieval latency from 10000 images (i.e., the database of Pitts30k) using L2 distance and the re-ranking latency of SelaVPR (100 candidates) are already within the same order of magnitude. When we reduce the dimension to 512, the retrieval latency also decreases by about 8 times. However, it is still 3.89ms, which is not negligible. In our new re-ranking paradigm, we use the Hamming distance to measure the similarity between binary features, which is more than an order of magnitude faster than calculating the L2 distance between floating-point features of the same dimension (only 0.26ms for 512-dim binary features). Overall, in terms of retrieval latency, while SelaVPR achieves more than an order of magnitude speedup over the common two-stage method, SelaVPR++ achieves more than two orders of magnitude speedup over SelaVPR and 6000× speedup over TransVPR. Compared to the common one-stage methods using 4096-dim global features, SelaVPR++ is also more than 60 times faster.
In addition, although we directly re-rank the top-100 candidates as in other works, it is also feasible to reduce the number of candidates (e.g., 50), which can further reduce latency. Table X shows the results of re-ranking different numbers of candidates. Our method achieves good results in all settings. There is no performance improvement when re-ranking top-200 candidates (compared to top-100), and a certain decrease when re-ranking top-20 candidates (compared to top-50). We believe that re-ranking 50-100 candidates is appropriate.
@@
TABLE VIII: The Results of Direct Retrieval Using Binary Features and Floating-Point Features, as Well as Our Re-Ranking Method. The R@100 Indicates the Upper Limit of (R@1/R@5/R@10) Performance After Re-Ranking, and Is Not Included in the Comparison.
@@
TABLE IX: The Retrieval Latency Comparison of Different Methods for a Single Query on the Pitts30k Dataset. The Initial Retrieval (to Get Candidates or Only One-Stage) is Denoted as “Initial”.
@@
TABLE X: The Results of Our Methods Using Different Numbers of Re-Ranking Candidates.
@@
4) Effect of our similarity-constrained deep hashing. Our deep hashing method combines the similarity-constrained loss and straight-through estimation (STE) [27]. The former is derived by simplifying the previous locality-constrained deep hashing work [70], which imposes a Gaussian prior on the pair-wise quantization loss. However, our work directly constrains the similarity of feature pairs before and after quantization without using the Gaussian prior. In this subsection, we compare the performance of direct hashing, locality-constrained (LC) hashing, STE, our similarity-constrained (SC) hashing, and our complete method (i.e., SC hashing + STE). To widen the performance gap between different methods, all models are based on the DINOv2-base backbone (outputting 512-dim features) and trained only on GSV-Cities. The results are shown in Table XI. We first summarize the results in terms of the average R@1/R@5 performance. The direct hashing method has the worst result because it uses the floating-point features when training the model and directly quantizes these features to obtain binary features during inference, which causes obvious information loss (or a gap between training and inference). LC hashing improves performance on all datasets by constraining the similarity of feature pairs before and after quantization, while our SC hashing simplifies LC hashing and achieves further improvement on the VPR task. Even so, STE achieves better overall performance than these methods by directly solving the gradient issue and using quantized binary features to calculate the metric loss for end-to-end training. Our complete method combines SC hashing and STE, and finally achieves the best performance among these methods. In addition, as for R@1 results on individual datasets, apart from our complete method, SC hashing achieves the best results on urban datasets (Pitts30k and Tokyo24/7) but is inferior to STE on suburban and natural datasets (MSLS and Nordland). Our approach yields outstanding results in different scenes simultaneously by combining these two methods.
@@
TABLE XI: The Results of Different Deep Hashing Methods. All Models Are Based on DINOv2-Base and Trained on the GSV-Cities Dataset.
@@
LC hashing [70] Pair-wise (Gaussian) $\times$ 88.4 94.8 87.8 94.5 87.0 94.9 55.2 74.7 79.6 89.7
SC hashing Pair-wise (simplified) $\times$ 89.1 94.7 88.0 93.8 89.2 95.9 57.1 75.6 80.9 90.0
@@
SC hashing + STE Pair-wise (simplified) 89.0 95.3 89.9 94.9 90.2 97.5 58.8 76.7 82.0 91.1
5) Effect of unifying and merging training datasets. The main contribution of the training strategy in our SelaVPR++ is to unify and merge several common VPR training datasets with the same training protocol for model training. To validate its effectiveness, we compare the performance of models trained with only GSV-Cities [16] and our proposed unified dataset. We conduct two sets of experiments that use both DINOv2-base and DINOv2-large backbone and produce 2048-dim and 4096-dim features respectively (as above). The results are shown in Table XII, and we summarize two points from it. (1) The model trained on the unified dataset consistently outperforms the one only using the GSV-Cities dataset on all evaluation datasets, regardless of whether DINOv2-base or DINOv2-large is used as the backbone. This demonstrates that training with our unified dataset can enhance the robustness and achieve better recognition performance. (2) The magnitude of performance improvement varies on different test datasets, with the largest gain achieved on Nordland (composed of natural landscapes), followed by MSLS-val (containing both urban and suburban scenes), and the smallest gains on Pitts30k and Tokyo24/7 (primarily consisting of urban street view images). For example, when using DINOv2-base, the R@1 improvements on Pitts30k, Tokyo24/7, MSLS-val, and Nordland are 0.7%, 0.6%, 1.7%, and 13.4%, respectively. This is because the GSV-Cities training set mainly consists of urban images, while our unified dataset covers more diverse VPR scenes. As a result, our model trained on the unified dataset is highly robust against diverse challenges, thus performing better and more consistently in various scenarios.
TABLE XII: The Results of Using the GSV-Cities and Unified Dataset for Training.

### v2

@@ IV-D Ablation Study
It is worth noting that the above fine-tuning methods are only applied to the backbone, while the aggregator (the GeM layer and the two linear layers before and after it) is always trainable. We use both DINOv2-base and DINOv2-large for experiments, outputting 2048-dim and 4096-dim global features for retrieval respectively. The GSV-Cities dataset is used for training. Since the Full-tuning will take up a huge amount of GPU memory, we adjust the batch size to 60 for all models based on DINOv2-large. Even so, the Full-tuning method still requires 6 NVIDIA 3090 GPUs. The results are shown in Table VI (DINOv2-base) and Table VII (DINOv2-large). We summarize the two tables in the following points: (1) For the Freezing method, since our aggregator contains two trainable linear layers that can also adapt the features produced by backbone to the VPR task, it achieves decent results for partial datasets. However, for the Nordland dataset, which is mainly collected in natural scenes and across seasons, the model needs to be able to capture detailed information in the image that can distinguish places. Simply adjusting the feature map output by the frozen backbone cannot effectively bridge the gap between the tasks of model pre-training and VPR. Therefore, the performance of this method is much lower than other methods. (2) As for Full-tuning, when using the DINOv2-base backbone, the difference between it and other methods is not large on most datasets. However, it is obviously inferior to other methods on Tokyo247, which contains many nighttime images. It is because there is a generalization gap between Tokyo24/7 and the GSV-Cities training dataset (without night images), and full fine-tuning damages the excellent representation ability of the pre-trained foundation model (i.e., catastrophic forgetting). This phenomenon is more obvious for DINOv2-large that has a larger number of parameters. (3) The performance of the PEFT methods (i.e., SelaVPR-global, MemEfficient-vanilla, and MemEfficient-MCA) when using DINOv2-large is generally better than that of DINOv2-base. However, the R@1 results of the Full-tuning and Partial-tuning methods on Tokyo247 decrease when using DINOv2-large. Especially for Full-tuning, the R@1 on Pitts30k and MSLS-val also decreases. This shows that the PEFT method is more suitable for large models than direct (full or partial) fine-tuning. (4) Both the Partial-tuning and PEFT methods achieve good performance on all these datasets. However, compared with the previously proven Partial-tuning and SelaVPR, the memory-efficient adaptation method used in this paper can achieve better results. Furthermore, our MemEfficent-MCA upgrades vanilla adapters to MultiConv adapters, which can facilitate feature interactions along the spatial axes and introduce the multi-scale local priors. Thus, it achieves the best performance among these methods.
@@
TABLE VI: The Results of Different Fine-Tuning Methods with the DINOv2-Base Backbone. Models Are Fine-Tuned on GSV-Cities and Batch Size Is Set to 120.
@@
TABLE VII: The Results of Different Fine-Tuning Methods with the DINOv2-Large Backbone. Models Are Fine-Tuned on GSV-Cities and Batch Size Is Set to 60.
@@
TABLE VIII: Comparison between the Memory-Efficient Adaptation Methods with the Vanilla Adapter, the Token-Channel Alternating Adapter, and Our MultiConv Adapter. We Use the DINOv2-Base Backbone, the GeM Aggregator (2048-Dim Output), and GSV-Cities for Training. The Resolution of the Input Image Is $224\times 224$ in Both Training and Inference.
Method Pitts30k-test MSLS-val Tokyo24/7 Nordland
R@1 R@5 R@1 R@5 R@1 R@5 R@1 R@5
MemEfficient-vanilla 91.2 95.7 90.3 95.6 93.7 96.5 69.2 83.2
MemEfficient-alternating 91.6 95.8 90.8 95.7 93.7 97.1 73.1 86.7
MemEfficient-MCA 91.8 96.2 91.2 95.7 94.0 97.5 75.2 87.9
@@
TABLE IX: The Comparison of Different Fine-Tuning Methods in Training Time for a Single Epoch on GSV-Cities. MemEfficient-MCA(4) Indicates Only Applying 4 Adapters for the Last 4 Blocks. We Consistently Train the Models on an NVIDIA A800 GPU with Batch Size Set to 40.
@@
Moreover, we also compare our method with the memory-efficient adaptation [29] that alternately applies adapters along the channel- and token-dimensions, denoted as MemEfficient-alternating. Due to the inherent limitation of the token-channel alternating adapter, it lacks adaptability to different image resolutions. Consequently, it can only be evaluated on images with the same resolution as in training. We compared three memory-efficient adaptation methods by setting the resolution to $\times224$ in both training and inference. The result is shown in Table VIII. Compared to MemEfficient-vanilla, MemEfficient-alternating enables patch token interactions, resulting in better results. Furthermore, our MultiConv adapter not only models spatial interactions between patch tokens but also introduces multiscale local priors beneficial for VPR, thus achieving the best results. More importantly, both MemEfficient-vanilla and ours can improve performance by increasing the resolution in inference (better results in Table VI using $\times322$ resolution in inference). However, this cannot be achieved for MemEfficient-alternating.
@@
Meanwhile, since the gradient backpropagation in our method no longer passes through the backbone during training, it saves a lot of training time. Table IX shows the training time of different fine-tuning methods for training an epoch on GSV-Cities (using an NVIDIA A800). Compared with Full-tuning, SelaVPR-global does not save significant time, that is, it is also not efficient in training time. For DINOv2-base, our MemEfficient-MCA consumes only 6.85 min/epoch, which is less than Partial-tuning (both are only slightly more than Freezing). For DINOv2-large, our method applies adapters to the last 16 blocks (DINOv2-base has only 12 blocks), while Partial-tuning still only fine-tunes the last 4 blocks. Therefore, the latter achieves a slightly shorter training time. However, our approach can further improve efficiency by reducing the number of adapters. When we use only 4 adapters, i.e., MemEfficient-MCA(4), the training time is 16.3% less than Partial-tuning. These results show that our method is efficient in training time in addition to parameter and memory. In the next ablation experiment, we will demonstrate that our method can still achieve excellent recognition performance when using only a small number of adapters.
TABLE X: The Results of Using Different Numbers of Adapters (with Two Ways). All Models Are Based on DINOv2-Large Backbone and Trained on Our Unified Dataset (Batch Size is 120).
@@
R@1 R@5 R@1 R@5 R@1 R@5 R@1 R@5 R@1 R@5
Freezing 92.7 96.8 92.0 96.8 95.6 98.4 69.7 82.8 87.5 93.7
All 94.3 97.4 94.7 97.7 97.5 98.7 96.2 98.8 95.7 98.2
Every 2 94.3 97.2 94.6 97.3 97.1 99.7 96.5 98.9 95.6 98.3
Every 3 94.4 97.5 94.2 97.3 97.5 98.7 96.4 98.9 95.6 98.1
Every 6 93.0 96.8 93.8 96.9 97.5 98.4 93.7 97.8 94.5 97.6
Last 16 94.4 97.5 94.3 98.0 98.1 99.0 97.0 99.1 96.0 98.4
Last 12 94.3 97.5 94.7 97.7 97.8 98.7 96.6 98.8 95.9 98.2
Last 8 94.0 96.8 94.9 97.6 97.1 99.7 95.5 98.6 95.4 98.2
Last 4 94.0 97.3 94.1 97.2 97.5 99.4 92.5 97.2 94.5 97.8
@@
2) Effect of the number of adapters. As described in Section III-B, our approach can further improve efficiency by reducing the number of adapters in two ways, namely, applying one adapter per $m$ blocks or applying adapters only to the last partial blocks. We equip DINOv2-large with different numbers of adapters and train them on our unified dataset, yielding the results shown in Table X. All methods equipped with adapters achieve good results and are significantly better than the baseline (i.e., Freezing). We mainly observe the performance of each setting through the average of R@N, which shows that the more adapters used in both ways, the better the performance. An exception is that the performance of applying adapters to the last 16 blocks is better than that for all blocks, indicating that the former has basically reached saturation performance. Therefore, for DINOv2-large, our recommendation is to only apply adapters to the last 16 blocks. In addition, the number of adapters used when applying an adapter per 2, 3, and 6 blocks just corresponds to applying adapters to the last 12, 8, and 4 blocks, and the results of the two are very close. So we believe both ways are good choices. Finally, reducing the number of adapters for the sake of efficiency is also advisable, because our method with only 4 adapters also achieves promising performance.
@@
TABLE XI: The Results of Direct Retrieval Using Binary Features and Floating-Point Features, as Well as Our Re-Ranking Method. The R@100 Indicates the Upper Limit of (R@1/R@5/R@10) Performance After Re-Ranking, and Is Not Included in the Comparison.
@@
TABLE XII: The Retrieval Latency Comparison of Different Methods for a Single Query on the Pitts30k Dataset. The Initial Retrieval (to Get Candidates or Only One-Stage) is Denoted as “Initial”.
@@
TABLE XIII: The Results of Our Methods Using Different Numbers of Re-Ranking Candidates.
@@
3) Effect of our novel two-stage paradigm. Our novel two-stage VPR paradigm discards the local features used in SelaVPR, and instead only uses global features for both initial retrieval and re-ranking. To demonstrate our re-ranking paradigm, we compare the performance of direct retrieval using binary features and floating-point features, as well as our re-ranking method (binary features for initial retrieval and floating-point features for re-ranking). We conduct experiments using both the DINOv2-base and DINOv2-large backbones with our unified dataset for training. The results are shown in Table XI. Since the binary features we use are also low-dimensional, their performance (R@1/R@5/R@10) is significantly lower than that of high-dimensional floating-point features. For the model using DINOv2-base, the R@1 margins on Pitts30k, MSLS-val, Tokyo247, and Nordland are 4.3%, 4.2%, 6.3%, and 16.3%, respectively. However, in our re-ranking pipeline, the initial search (using binary features) is only expected to return the top-100 candidate images with correct results. The difference between the R@100 results obtained by binary and floating-point features is very slight. After re-ranking the top-100 candidates using high-dimensional floating-point features, the R@1/R@5/R@10 results of our method are not significantly different from those yielded by directly using high-dimensional floating-point features (some are slightly higher, and some are slightly lower).
Since our method uses compact binary features to provide candidate images by calculating Hamming distances, this can further greatly improve retrieval efficiency. Table XII shows the retrieval latency (including initial retrieval latency and re-ranking latency) of a single query on Pitts30k for three two-stage methods, i.e., TransVPR [12], SelaVPR [13], and SelaVPR++. We also provide the latency of direct retrieval with the L2 distance for 4096-dim, 2048-dim, and 512-dim floating-point features, as well as with the Hamming distance for 512-dim binary features. Although our previous work SelaVPR matches local features for re-ranking, it eliminates the time-consuming geometric verification process in the common two-stage method (e.g., TransVPR). Its re-ranking latency is thus more than one order of magnitude lower than that of TransVPR. Our SelaVPR++ directly uses global features for re-ranking, which is further more than two orders of magnitude faster than SelaVPR. As for the initial retrieval latency, once the measurement is determined (e.g., L2 distance), the latency is roughly proportional to the feature dimension. When the feature dimension is 4096, the retrieval latency from 10000 images (i.e., the database of Pitts30k) using L2 distance and the re-ranking latency of SelaVPR (100 candidates) are already within the same order of magnitude. When we reduce the dimension to 512, the retrieval latency also decreases by about 8 times. However, it is still 3.89ms, which is not negligible. In our new re-ranking paradigm, we use the Hamming distance to measure the similarity between binary features, which is more than an order of magnitude faster than calculating the L2 distance between floating-point features of the same dimension (only 0.26ms for 512-dim binary features). Overall, in terms of retrieval latency, while SelaVPR achieves more than an order of magnitude speedup over the common two-stage method, SelaVPR++ achieves more than two orders of magnitude speedup over SelaVPR and 6000× speedup over TransVPR. Compared to the common one-stage methods using 4096-dim global features, SelaVPR++ is also more than 60 times faster.
In addition, although we directly re-rank the top-100 candidates as in other works, it is also feasible to reduce the number of candidates (e.g., 50), which can further reduce latency. Table XIII shows the results of re-ranking different numbers of candidates. Our method achieves good results in all settings. There is no performance improvement when re-ranking top-200 candidates (compared to top-100), and a certain decrease when re-ranking top-20 candidates (compared to top-50). We believe that re-ranking 50-100 candidates is appropriate.
@@
TABLE XIV: The Results of Different Deep Hashing Methods. All Models Are Based on DINOv2-Base and Trained on the GSV-Cities Dataset.
@@
LC hashing [74] Pair-wise similarity [74] $\times$ 88.4 94.8 87.8 94.5 87.0 94.9 55.2 74.7 79.6 89.7
SC hashing Pair-wise similarity (ours) $\times$ 89.1 94.7 88.0 93.8 89.2 95.9 57.1 75.6 80.9 90.0
@@
SC hashing + STE Pair-wise similarity (ours) 89.0 95.3 89.9 94.9 90.2 97.5 58.8 76.7 82.0 91.1
TABLE XV: The Results of Our Method Using the GSV-Cities and Unified Dataset for Training.
@@
4) Effect of our similarity-constrained deep hashing. Our deep hashing method combines the similarity-constrained loss and straight-through estimation (STE) [28]. The former is inspired by the locality-constrained deep hashing [74], which preserves pair-wise similarity (inner product) between images and uses a sigmoid transformation on the inner product. However, our work directly constrains the cosine similarity of feature pairs before and after quantization without using the sigmoid function. In this subsection, we compare the performance of direct hashing, locality-constrained (LC) hashing, STE, our similarity-constrained (SC) hashing, and our complete method (i.e., SC hashing + STE). To widen the performance gap between different methods, all models are based on the DINOv2-base backbone (outputting 512-dim features) and trained only on GSV-Cities. The results are shown in Table XIV. We first summarize the results in terms of the average R@1/R@5 performance. The direct hashing method has the worst result because it uses the floating-point features when training the model and directly quantizes these features to obtain binary features during inference, which causes obvious information loss (or a gap between training and inference). LC hashing improves performance on all datasets by constraining the similarity of feature pairs before and after quantization, while our SC hashing is not only more concise than LC hashing but also achieves further improvement on the VPR task. Even so, STE achieves better overall performance than these methods by directly solving the gradient issue and using quantized binary features to calculate the metric loss for end-to-end training. Our complete method combines SC hashing and STE, and finally achieves the best performance among these methods. In addition, as for R@1 results on individual datasets, apart from our complete method, SC hashing achieves the best results on urban datasets (Pitts30k and Tokyo24/7) but is inferior to STE on suburban and natural datasets (MSLS and Nordland). Our approach yields outstanding results in different scenes simultaneously by combining these two methods.
5) Effect of unifying and merging training datasets. The main contribution of the training strategy in our SelaVPR++ is to unify and merge several common VPR training datasets with the same training protocol for model training. To validate its effectiveness, we compare the performance of models trained with only GSV-Cities [16] and our proposed unified dataset. We conduct two sets of experiments that use both DINOv2-base and DINOv2-large backbones and produce 2048-dim and 4096-dim features respectively (as above). The results are shown in Table XV, and we summarize two points from it. (1) The model trained on the unified dataset consistently outperforms the one only using the GSV-Cities dataset on all evaluation datasets, regardless of whether DINOv2-base or DINOv2-large is used as the backbone. This demonstrates that training with our unified dataset can enhance the robustness and achieve better recognition performance. (2) The magnitude of performance improvement varies on different test datasets, with the largest gain achieved on Nordland (composed of natural landscapes), followed by MSLS-val (containing both urban and suburban scenes), and the smallest gains on Pitts30k and Tokyo24/7 (primarily consisting of urban street view images). For example, when using DINOv2-base, the R@1 improvements on Pitts30k, Tokyo24/7, MSLS-val, and Nordland are 0.7%, 0.6%, 1.7%, and 13.4%, respectively. This is because the GSV-Cities training set mainly consists of urban images, while our unified dataset covers more diverse VPR scenes. As a result, our model trained on the unified dataset is highly robust against diverse challenges, thus performing better and more consistently in various scenarios.
Figure 8: Some examples of corner cases (success cases in the first three and failure cases in the last two). For the first example, the query is rotated 180. All other methods return the inverted but wrong results, while only our SelaVPR++ gets the correct image (with normal orientation). The second example shows a nighttime query with weak light and limited visibility. Our method is highly robust against light changes and gets the right database image (exhibiting obvious viewpoint changes with the query), but others fail. The third and fourth query images are taken in nature, lacking discriminative landmarks and buildings. Our method, which can capture other valuable details (e.g., terrain characteristics and vegetation), succeeds in the third case. However, this does not guarantee a correct match every time, especially when the plants shown in the query image lack distinctive characteristics and occupy a majority of the image region (i.e., the fourth case). In the last example, most methods return the same result that is geographically close to the query but exceeds the predefined threshold (i.e., failed). Although SALAD-CM introduces a novel CliqueMining strategy explicitly tailored to address this issue, it still yields the same wrong result as ours.